OpenStudy (anonymous):

Find the sum of each of the geometric series given below. For the value of the sum, find an expression that gives the exact value, rather than giving an approximation. 1) -6 + 2 - (2/3) + (2/9) - (2/27) + (2/81) - ... 2)(Sigma from n=5 to n=12) (1/3)^n

OpenStudy (zzr0ck3r):

bah this is to hard on my tablet, brb

OpenStudy (zzr0ck3r):

ok the first term is -6 and the common ratio is -1/3 so we have -6/(1-(-1/3)) = -9/2

OpenStudy (zzr0ck3r):

\(\sum_{0}^\infty ar^n=\frac{a}{1-r}\) as long as \(|r|<1\)

OpenStudy (anonymous):

With r being the ratio.

OpenStudy (zzr0ck3r):

your first one is \(\sum_0^\infty(-1)^{n+1}6(\frac{1}{3})^n\)

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

what is the ratio of the second one?

OpenStudy (anonymous):

Hold on. Why not make the first one similar, but with this as the equation instead: -6(-1/3)^(n-1)

OpenStudy (anonymous):

Because with the first number in the series, that would make it positive two, instead of negative 6, correct?

hartnn (hartnn):

you can do that. to find the sum of infinite geometric series, we need 2 things: 1) first term 2) common ratio no matter how you find it. so whats first term and ratio for 2nd part ?

OpenStudy (anonymous):

First term would be 1/3, with a ratio of 1/3?

OpenStudy (anonymous):

For the second problem, correct?

hartnn (hartnn):

oops, i din't note that 2nd problem is not of infinite sequence its finite, n= 5 to 12

OpenStudy (anonymous):

Yeah, I was a bit confused there! Lol, I ended up getting the first one, but this second one is giving me problems.

hartnn (hartnn):

the formula you use is : \(\large S_n=a_1\frac{(1-r^n)}{(1-r)}\) a1 = 1st term , n= number of terms r = ratio = 1/2 to get first term, put n=5 in (1/3)^n because n starts from 5

hartnn (hartnn):

to get number of terms, well, from 5 to 12, there are 12-5+1 terms so your n in the formula will be n =12-5+1

OpenStudy (anonymous):

Six terms. Gotcha.

OpenStudy (anonymous):

Ended up not getting it quite right. I'll try to work on it tomorrow, letting this thread sit overnight.

hartnn (hartnn):

12-5 is 7 and 7+1 is 8 5,6,7,8,9,10,11,12 n=8 a1 = (1/3)^5 r =1/3 :)

OpenStudy (anonymous):

Bah! I feel dumb now. Did my addition and subtraction wrong. D: Thank you!